# Enhancing Vision-Language Models with Scene Graphs for Traffic Accident Understanding

### v1

@@ IV-D Implementation Details
We train the SGE for 200 epochs on the default hyperparameter settings as specified by Yu et al. (2020). For the SGE pre-training experiments, we select the SGE with the lowest validation loss. We also experiment with the provided SGE pre-trained checkpoint that was trained on synthetic traffic scenes generated by CARLA. For the modality alignment training, we keep most of the default hyperparameter settings, varying batch size (32 to 128) and the number of epochs (8 to 20).
@@ V-A SGE Pre-training Results
As seen in Table I, pre-training the SGE on the DoTA dataset allows it to perform the classification task approximately 13 percentage points better than random classification, which would have an expected accuracy of 25%. This suggests that the spatial and temporal information encoded in the graph embeddings produced by the SGE contain useful information towards accident classification. Unsurprisingly, the model pre-trained on the synthetic CARLA dataset does not achieve as strong performance on our task as the model trained on DoTA. This is understandable, given that the former was trained to classify whether or not a traffic maneuver is deemed to be risky Malawade et al. (2022) rather than to classify traffic accidents. Furthermore, the synthetic traffic scenes in the CARLA dataset used by the authors do not vary as much as those seen in DoTA, and do not contain examples from the four classes used for our task. However, by further training the CARLA model on the DoTA dataset, we are able to achieve similar results to training on DoTA alone (compare two rightmost columns of Table I). For the remaining experiments in the paper, when unspecified, we use the SGE pre-trained only on the DoTA dataset.
@@
Accuracy
Train - 62.72 64.09
Train (b) - 64.78 61.42
Validation - 42.49 41.91
Validation (b) - 40.79 41.42
Test 17.32 37.88 36.03
Test (b) 26.33 38.05 37.81
TABLE II: Test Accuracy Comparison between three system settings. In the first setting, the SGE is not used and the inputs to the system are only from the video and text modalities (the SGG and SGE are unused and stage III-C from the pipeline is skipped). In the second, the SGE is used and aligned with the vision and text encoders before running classification (full architecture is used, but SGE pre-training done in stage III-B may be skipped). In the third, the SGE is used but not aligned with the other encoders (stage III-C is skipped).
@@ V-B SGE with Alignment Results
The primary classification results after training for alignment can be seen in Table II. In Column 1, we verify the baseline accuracy of the model when the scene graph modality is not used at all. Looking at the balanced accuracy score, it is quite interesting to note that the performance under this setting is very close to that of the aligned model when no pre-training data is used or if only the CARLA dataset is used (Columns 2, 3). However, pre-training the model on the DoTA dataset (Column 4) shows a boost in performance by about 3%. This boost suggests that the SG embeddings may be encoding some information about the traffic scenes that cannot be fully captured by either the language or vision modalities.
@@ V-B1 Pre-trained SGE with Alignment
Regarding the case when the model is pre-trained on the DoTA dataset, the model actually has a slightly better performance without undergoing the alignment training (Column 5) versus with alignment (Column 4). This seems to suggest that the model is not learning further useful information during alignment training. One possible explanation for this is that the data domains of the frozen encoders are not similar enough to DoTA for any positive transfer of data domain knowledge to occur Zhang et al. (2023). Another possibility has to do with the alignment training procedure, which is based on CLIP’s training algorithm and relies heavily on using a large batch size at each step Radford et al. (2021). This is because the algorithm performs contrastive training on each of the three pairs of embedding spaces (vision-graph, text-graph, and vision-text) for each batch of traffic accident examples. However, given the limited size of the DoTA dataset and compute constraints, our experiments in Table II use a batch size of 32 and are limited to 10 epochs of training. With a larger batch size (and possibly increasing the number of training epochs), it is possible that the hypothesized trend of higher accuracy when training for alignment would be present. To this end, we show some additional results in Table III, which indicate that both increasing batch size and the number of epochs give similar results to the “SGE without Alignment” case shown in Table II (Column 5) rather than worse results (Columns 2-4 of Table II). It is possible that further training of the larger batch sizes would have improved these results beyond the values in Column 5.
TABLE III: Pre-training on DoTA with different configurations. “e” is number of epochs, “bs” is batch size. Column 4 in Table II is the same as Column 1 in this table.
@@ V-B2 No Pre-trained SGE with Alignment
Another surprising point that arises from the results presented in Table II is that there is very little improvement when training the SGE for alignment from scratch (Column 2) over the baseline “No SGE” case (Column 1). Similar to the case where we pre-trained on DoTA, it is once again possible that further increasing the batch size could be the solution. This is suggested by the results in Table IV, where larger batch sizes begin to show a significant increase in performance over the baseline.
@@ VI Conclusion and Future Work
In this work, we have shown that encoding traffic information in the form of a scene graph is beneficial towards the goal of accident classification. This is made clear by the pre-training encoder results which show the SGE’s ability to beat a random classifier at this task. It was further illustrated that the scene graph information can serve to enhance the performance of a vision-language classifier by fusing information from all three modalities. Finally, although we were not able to show an improved score for the classifier after alignment, it should be noted that alignment does not have a negative effect on performance, and we demonstrate that further training and fine-tuning may improve the score beyond the unaligned case.
There are several areas where future work can be done in this task. Firstly, although three modalities are used, little exploration has been conducted on the language modality. In principle, captions for each video should be derived from the video itself, as the SGs are. This would allow for the captions to be fed into the model to fine-tune the classifier head and run inference with a greater signal coming from the language modality. Next, although they provide some benefit for the model, the generated SGs have a relatively limited structure: only categorical distance relations between a vehicle and objects in its surroundings are generated, along with mappings to a fixed set of three traffic lanes (left, middle, and right). Enabling a semantic extension of the generated SGs may enhance the signal provided from this modality. Similarly, by modeling relations between all objects in the scene, this pipeline can be expanded for the non-ego case as well, and perhaps can be used on more classes in the DoTA dataset or similar use cases. Additionally, the MRGCN-based architecture of the SGE was taken directly from Yu et al. (2020), but perhaps it can be further improved by modifying either its spatial or temporal modeling components. This work has demonstrated that further fine-tuning and alignment training shows promise to improve results. Experiments with different modality fusion methods as well as classification heads can also lead to improvements.
Appendix A Generated Scene Graph Example

### v2

@@ IV-D Implementation Details
We train the SGE for 200 epochs on the default hyperparameter settings as specified by Yu et al. (2020). After the SGE pre-training experiments, we select the SGE with the lowest validation loss to use for the modality alignment training. During alignment training, we keep most of the default hyperparameter settings, except for varying the batch size (32 to 128) and the number of epochs (8 to 20).
@@ V-A SGE Pre-training Results
As seen in Table I, pre-training the SGE on DoTA allows it to perform classification approximately 13 percentage points better than random classification, which would have an expected accuracy of 25%. This suggests that spatial and temporal information encoded in the graph embeddings produced by the SGE contain useful information towards accident classification. Unsurprisingly, the model pre-trained on the synthetic CARLA dataset does not achieve as strong performance on our task as the model trained on DoTA. This is understandable, given that the former was trained to classify whether or not a traffic maneuver is deemed to be risky Malawade et al. (2022) rather than to classify traffic accidents. Furthermore, the scenes in CARLA do not vary as much as those in DoTA, and do not contain examples from the four classes we use. However, by further training the CARLA model on DoTA, we achieve similar results to training on DoTA alone (compare two rightmost columns of Table I). For the remaining experiments in the paper, when unspecified, we use the SGE pre-trained only on the DoTA dataset.
@@
Train accuracy - 62.72 64.09
Train accuracy (b) - 64.78 61.42
Validation accuracy - 42.49 41.91
Validation accuracy (b) - 40.79 41.42
Test accuracy 17.32 37.88 36.03
Test accuracy (b) 26.33 38.05 37.81
TABLE II: Test Accuracy Comparison between three system settings. In the first setting, the SGE is not used and the inputs to the system are only from the video and text modalities (the SGG and SGE are unused and stage III-C from the pipeline is skipped). In the second, the SGE is used and aligned with the vision and text encoders before running classification (full architecture is used, but SGE pre-training done in stage III-B may be skipped). In the third, the SGE is used but not aligned with the other encoders (stage III-C is skipped). All values are obtained using a batch size of 32 and 10 epochs of training.
@@ V-B SGE with Alignment Results
The primary classification results after training for alignment can be seen in Table II. In Column 1, we verify the baseline accuracy of the model when the scene graph modality is not used at all. Looking at the balanced accuracy score, it is interesting to note that the performance on this setting is only slightly under that of the aligned model when no pre-training data is used or if only the CARLA dataset is used (Columns 2, 3). However, pre-training the model on the DoTA dataset (Column 4) shows a further boost in performance by about 3%. This boost suggests that the SG embeddings may be encoding some information about the traffic scenes that cannot be fully captured by either the language or vision modalities.
@@ V-B1 Pre-trained SGE with Alignment
Regarding the case when the model is pre-trained on the DoTA dataset, the model actually has a slightly better performance without undergoing the alignment training (Column 5) versus with alignment (Column 4). This seems to suggest that the model is not learning further useful information during alignment training. One possible explanation for this is that the data domains of the frozen encoders are not similar enough to DoTA for any positive transfer of data domain knowledge to occur Zhang et al. (2023). Another possibility has to do with the alignment training procedure, which is based on CLIP’s training algorithm and therefore relies heavily on using a large batch size at each step Radford et al. (2021). This is because the algorithm performs contrastive training on each of the three pairs of embedding spaces (vision-graph, text-graph, and vision-text) for each batch of traffic accident examples. However, given the limited size of the DoTA dataset and compute constraints, our experiments in Table II use a batch size of 32 and are limited to 10 epochs of training. We hypothesize that a larger batch size and/or number of training epochs would increase the accuracy gain of the training alignment step. To test this hypothesis, we carried out some additional experiments, the results of which are shown in Table III, and indicate that both increasing batch size and the number of epochs give similar results to the “SGE without Alignment” case shown in Table II (Column 5) rather than worse results (Columns 2-4 of Table II). It is possible that further training of the larger batch sizes would have improved these results beyond the values in Column 5.
TABLE III: Alignment training with different hyperparameter configurations, using a SGE pre-trained on DoTA. “e” is number of epochs, “bs” is batch size for the alignment training. Column 4 in Table II is the same as Column 1 in this table.
@@ V-B2 No Pre-trained SGE with Alignment
Another surprising point arising from Table II, which was mentioned at the start of Section V-B, is that there is very little improvement when training the SGE for alignment from scratch (Column 2) over the baseline “No SGE” case (Column 1). Similar to the case where we pre-trained on DoTA, it is once again possible that further increasing the batch size could be the solution. This is corroborated by the results in Table IV, where larger batch sizes begin to show a significant increase in performance over the baseline.
@@ VI Conclusion and Future Work
In this work, we have shown that encoding traffic information in the form of a scene graph is beneficial towards the goal of accident classification. This is made clear by the pre-training encoder results which show the SGE’s ability to beat a random classifier at this task. It was further illustrated that scene graph information can enhance the performance of a vision-language classifier by fusing information from all three modalities. Finally, although we were not able to show an improved score for the classifier after alignment, it should be noted that alignment training using an appropriate hyperparameter configuration does not have a negative effect on performance, and we demonstrated that further training and fine-tuning may improve the score beyond the unaligned case.
There are several areas where future work can be done in this task. Firstly, although three modalities are used, little exploration has been conducted on the language modality. In principle, captions for each video should be derived from the video itself, as the SGs are. This would allow for the captions to be fed into the model to fine-tune the classifier head and run inference with a greater signal coming from the language modality. Next, although they provide some benefit for the model, the generated SGs have a relatively limited structure: only categorical distance relations between a vehicle and objects in its surroundings are generated, along with mappings to a fixed set of three traffic lanes (left, middle, and right). Enabling a semantic extension of the generated SGs may enhance the signal obtained from this modality. Similarly, by modeling relations between all objects in the scene, this pipeline could be expanded for the non-ego case as well, and perhaps could be used on more accident classes in the DoTA dataset or similar use cases. Regarding the dataset, although we attempt to resolve the data imbalance by focusing on the balanced accuracy metric, it would be interesting to explore data augmentation techniques as well. Additionally, the MRGCN-based architecture of the SGE was taken directly from Yu et al. (2020), but perhaps it can be further improved by modifying either its spatial or temporal modeling components. Given the modular design of our system, another fascinating avenue of exploration would be to incorporate an additional modality (e.g., audio) to include in the alignment step before fusing the embeddings and fine tuning the classification head. Experiments with different modality fusion methods as well as classification heads may also lead to improvements.
Appendix A Generated Scene Graph Examples
@@
Figure 3: Starting from a video frame, the SGG first detects objects in the scene and generates the BEV image before creating the scene graph representation. This scene graph shows the ego car relative to a person on a scooter, categorized as in the left lane. The person is recognized as being near collision, with the edge attribute “near_coll”. Note that there is a vehicle detected in the distance (right part of detection image, highlighted in blue) but it is not included in the graph since it is too far.